OpenStudy (anonymous):

¡WILL MEDAL! Simplify the following expression. (67x + 66y) + (-12x + 50y) A.55x + 16y B.79x + 16y C.79x + 116y D.55x + 116y

OpenStudy (anonymous):

u have to add the like terms

OpenStudy (anonymous):

Add like terms. _x + _x = _x _y + _y = _y

OpenStudy (anonymous):

Let's simplify step-by-step. 67x+66y−12x+50y =67x+66y+−12x+50y Combine Like Terms: =67x+66y+−12x+50y =(67x+−12x)+(66y+50y) =55x+116y Answer: =55x+116y

OpenStudy (anonymous):

@byronichero